# Simple Semi-supervised Knowledge Distillation from Vision-Language Models via Dual-Head Optimization

### v1

@@ D.1 Experiments on MobileNet
Figure 11: Results on 10 datasets under few-shot semi-supervision using MobileNetV2 with zero-shot teacher [72].
@@
Figure 12: Results on 10 datasets using MobileNetV2 with either zero- or few-shot teacher [115].
@@
To demonstrate the versatility of our DHO approach beyond ResNet [32] and ViT [20] architectures, we extended our experiments to the MobileNetV2 [77], which is specifically designed for real-world applications with compact models. We maintained identical experimental settings as described in Section B.2, using MobileNetV2 as the student model while distilling from CLIP ResNet50.
@@
As illustrated in Figure 12, our DHO consistently outperforms all single-head baseline methods, demonstrating its effectiveness on lightweight model architectures along with ResNet18. Furthermore, Figure 12 reveals patterns similar to our ResNet18 experiments regarding few-shot integration. Our method successfully incorporates few-shot teacher knowledge, although we observe that the few-shot teacher does not consistently yield improvements over the zero-shot teacher. Notably, our distilled MobileNetV2 model sometimes achieves superior performance to the few-shot teacher (ResNet-50) despite having significantly fewer parameters.
D.2 Out-of-Distribution Generalization
While our DHO framework is primarily designed for task-specific adaptation, we extensively evaluate its generalization capabilities across distribution-shifted scenarios. We conduct comprehensive experiments on four widely-used distribution-shifted variants of ImageNet: ImageNet-v2 [73], ImageNet-Sketch [93], ImageNet-R [36], and ImageNet-A [37] (detailed dataset characteristics are presented in Table 10).
Full Training
Model Configuration Training Details
Student: CLIP ViT-B/16 [72] Student input size: 224$\times$224 Zero-shot Teacher: CLIP ViT-L/14 [72] Teacher input size: 336$\times$336 Labeled data: 1% and 10% ImageNet $\zeta$, $\eta$, and $\lambda$: $0.01,2,0.5$ $\alpha$ and $\beta$: $0.5$ and $1$ Epochs: 32 Optimizer: AdamW ($\beta_{1}$=0.9, $\beta_{2}$=0.999) Learning rate: $5\times 10^{-5}$, weight decay: $5\times 10^{-2}$ Batch size: 512 (labeled: 256, unlabeled: 256) Scheduler: Cosine warmup decay (5000 steps) Augmentation: Random crops (x0.5-1.0), horizontal flips
Adaptation Methods (Linear Evaluation & Visual Prompt Tuning)
Method Configuration Training Details
Linear evaluation [8] Visual prompt tuning [43] Frozen backbone: CLIP ViT-B/16 [72] Input size: 224$\times$224 Zero-shot Teacher: CLIP ViT-L/14 [72] Teacher input size: 336$\times$336 Labeled data: 1% and 10% ImageNet $\zeta$, $\eta$, and $\lambda$: $0.01,2,0.5$ $\alpha$ and $\beta$: $0.5$ and $1$ Epochs: 20 Optimizer: AdamW ($\beta_{1}$=0.9, $\beta_{2}$=0.999) Learning rate: $5\times 10^{-5}$, weight decay: $5\times 10^{-2}$ Batch size: 512 (labeled: 256, unlabeled: 256) Scheduler: Cosine warmup decay (5000 steps) Augmentation: Random crops (x0.5-1.0), horizontal flips
Adaptation Methods (Prompt Tuning)
Method Configuration Training Details
Prompt tuning: CoOp [120], PromptSRC [45] Frozen backbone: CLIP ViT-B/16 [72] Input size: 224$\times$224 Zero-shot Teacher: CLIP ViT-L/14 [72] Teacher input size: 336$\times$336 Labeled data: 1% and 10% ImageNet $\zeta$, $\eta$, and $\lambda$: $0.01,2,0.5$ $\alpha$ and $\beta$: $0.5$ and $1$ Prompt tuning: Following PromptSRC [45] configurations Comparison: CasPL [97] with domain-specific unlabeled data
Our analysis consists of two main components. First, we directly evaluate our ImageNet semi-supervised models on these distribution-shifted benchmarks to assess how DHO impacts robustness. Second, we investigate the integration of established adaptation techniques while preserving generalization capabilities. Our investigation begins with linear evaluation protocols [72] and Visual Prompt Tuning [43] using frozen image backbones. We then extend our analysis to state-of-the-art CLIP-based few-shot adaptation methods, specifically examining CoOp [120] for unimodal prompt tuning and PromptSRC [45] for multimodal prompt tuning approaches. Additionally, we compare our framework with CasPL [97], which leverages domain-specific unlabeled data with separate distillation stage to enhance adaptation performance.
To evaluate our approach’s effectiveness under different data availability scenarios, we conducted experiments under both 1% and 10% labeled data settings, maintaining balanced loss weights of 0.5-0.5 and fixed values of $\alpha=0.5$ and $\beta=1.0$ for inference. Throughout the training process, we ensured consistent conditions by maintaining a 1:1 ratio of labeled to unlabeled samples in the batch composition. For experiments utilizing only labeled samples, we matched the number of training steps to those experiments using both labeled and unlabeled data to ensure fair comparison.
Evaluation upon Full Training Model.
@@ D.2 Out-of-Distribution Generalization
ViT-H [22] 632M zero-shot - 83.6 77.2 71.7 92.3 77.4
DHO significantly outperformed zero-shot baselines on similar-distribution ImageNet variants (Table 12), demonstrating particularly strong performance on ImageNet-V2 while maintaining reasonable accuracy on ImageNet-Sketch. This performance pattern remained consistent across both ViT-B/16 and ViT-L/14 architectures, with our approach showing improvements on similar-distribution test sets. However, when evaluated on out-of-distribution datasets such as ImageNet-R and ImageNet-A, the model exhibited performance degradation compared to zero-shot predictions, suggesting that full model training led to increased distribution overfitting and compromised generalization capability across domains.
A particularly insightful observation emerged from our analysis of different teacher-student configurations. Interestingly, ViT-B/16 models distilled from ViT-L/14 handled shifted distributions better than those taught by the larger ViT-H/14 DFN [22], despite the latter’s superior performance on shifted distributions such as ImageNet-R and ImageNet-A. We attribute this to the shared training background between ViT-B/16 and ViT-L/14 in the CLIP framework [72], which appears to better preserve generalization capabilities during the adaptation. This points to an important insight: our method works best on out-of-distributions when the teacher and student models share similar training distributions, suggesting that successful knowledge distillation also depends on the alignment between teacher and student than just the teacher’s raw capabilities.
Evaluation upon Adaptation Methods.
Table 13: Accuracy(%) of DHO with adaptation methods on the ImageNet distribution-shifted variants. $\dagger$: backbone frozen during training. VPT-$n$ indicates Visual Prompt Tuning with prompt depth of $n$ layers. For fair comparison, we used $\alpha=0.5$ and $\beta=1.0$ for all DHO evaluation.
Method Labeled Teacher Unlabeled Val V2 Sketch R A
Data Model Data
CLIP Zero-shot
CLIP [72] - - - 66.7 60.8 46.2 74.0 47.8
Linear Evaluation
Linear Evaluation 1% - - 72.8 64.2 47.0 74.6 48.6
DHO 1% - 73.4 65.3 48.3 75.8 49.4
DHO 1% 74.6 66.4 49.2 76.1 49.8
Visual Prompt Tuning
VPT-3 [43] 1.25% - - 73.6 64.6 47.7 75.2 48.7
VPT-3+DHO 1.25% - 73.6 65.3 48.7 75.9 49.1
VPT-3+DHO 1.25% 75.1 66.8 50.0 76.9 50.5
VPT-9+DHO 1.25% - 74.0 65.9 48.9 75.8 46.5
VPT-9+DHO 1.25% 75.6 67.4 50.8 77.3 46.5
VLM Text-encoder Prompt Tuning
CoOp [120] 1.25% (16-shots) - - 71.5 64.2 48.0 75.2 49.7
CoOp+CasPL [97] 1.25% (16-shots) 71.9 64.3 48.3 76.0 -
CoOp+DHO 1.25% (16-shots) - 72.8 65.5 49.3 76.4 49.5
CoOp+DHO 1.25% (16-shots) 73.4 66.2 49.5 77.0 50.5
VLM Multimodal Prompt Tuning
PromptSRC [45] 1.25% (16-shots) - - 71.3 64.4 49.6 77.8 50.9
PromptSRC+CasPL [97] 1.25% (16-shots) 72.8 65.7 49.7 77.9 -
PromptSRC+DHO 1.25% (16-shots) - 73.0 65.3 49.5 77.8 51.3
PromptSRC+DHO 1.25% (16-shots) 73.6 66.1 49.8 78.1 51.0
As demonstrated in Table 13, we first evaluated DHO in scenarios where only the image encoder is utilized: linear evaluation [72] and visual prompt tuning [43] with a frozen CLIP ViT-B/16 encoder. Notably, we discovered that linear evaluation with language initialization, as proposed in [51], performs remarkably well, even outperforming prompt tuning-based methods such as CoOp [120] and PromptSRC [45] on in-domain distribution benchmarks. With the additional pathway for teacher knowledge distillation, DHO achieves substantial performance improvements across all in-distribution and out-of-distribution evaluations, even with just 1% of labeled data. Furthermore, leveraging unlabeled data yields additional performance gains.
In experiments with visual prompt tuning, similar to our linear probing experiments, introducing a new optimization pathway for distillation resulted in performance improvements across all evaluation benchmarks, with further enhancements when utilizing unlabeled data. Moreover, increasing the number of learnable parameters by adding prompt layers led to improved performance across all datasets, with ImageNet-A being the sole exception.
We also evaluated DHO on various VLM adaptation methods, including the text-encoder prompt learning approach CoOp [120], the multi-modal prompt learning method PromptSRC [45], and the recently proposed CasPL [97], which employs a teacher network for unsupervised distillation on unlabeled data as a separate stage. We used the same implementation from the domain generalization setting of the original papers [120, 45], with only a linear layer jointly trained on top of image encoder features as the distillation pathway.
Our results demonstrate that the simple integration of DHO with existing approaches yields remarkable outcomes as shown in Table 13, achieving average improvements of 0.9% and 0.3% across four generalization benchmarks upon CoOp and PromptSRC, respectively. Notably, when compared to CasPL, which utilizes a teacher network with domain unlabeled datasets, our DHO approach demonstrates superior performance across all in-distribution and out-of-distribution benchmarks, achieving new state-of-the-art results when leveraging both teacher and unlabeled domain datasets. This suggests that joint training of labeled supervision and teacher distillation signals is a more effective strategy than the sequential approach used in CasPL, aligning with our core hypothesis about the benefits of dual-head optimization for resolving gradient conflicts.
Appendix E Additional Analysis
E.1 Gradient Analysis
Figure 13: (Left): The average dominant eigenvalue $\lambda$ of $WW^{\top}$ and $W_{\text{CE}}W_{\text{KD}}^{\top}$ during training. (Right): The average cosine similarity of gradients w.r.t features $\mathtt{CosSim}(\nabla_{z}\mathcal{L}_{\text{CE}},\nabla_{z}\mathcal{L}_{\text{KD}})$.
We analyze the gradient conflict occurs in one of single-head KD baselines, logit distillation in §3.2. In this section, we present the dominant eigen value of $WW^{\top}$ and $W_{\text{CE}}W_{\text{KD}}$ during training in Figure 13-(Left). We further present the cosine similarity between gradients w.r.t feature representation $z$ in Figure 13-(Right). The trend of cosine similarity between feature gradients is similar to that of full gradients $\mathtt{CosSim}(\nabla_{\theta}\mathcal{L}_{\text{CE}},\nabla_{\theta}\mathcal{L}_{\text{KD}})$ in Figure 4.
E.2 Additional Analysis on Non-Linear Head Design
@@
To further investigate the architectural advantages of dual head optimization, we conducted experiments with non-linear head designs, replacing the linear heads used in our main experiments. The non-linear classifier enhanced model representational capacity through a sequence of layers: an initial linear projection layer, followed by layer normalization [4], GELU activation [35], dropout [82], and a final linear classification layer. We compared DHO with three non-linear configurations; DHO+NL-Head-CE: non-linear CE head, DHO+NL-Head-KD: non-linear KD head, and DHO+NL-Head-CE+KD: non-linear both CE and KD heads. All experiments followed the few-shot semi-supervised setting detailed in Section B.2.
Figure 14: Results of different non-linear head configurations on 11 datasets including ImageNet under few-shot semi-supervision using ResNet-18 with zero-shot teacher [72].
@@
Table 14: Results of different non-linear head configurations on 11 datasets including ImageNet under few-shot semi-supervision using ResNet-18 with zero-shot teacher [72]. We report averaged accuracy for 10 visual recognition datasets except the ImageNet.
@@
Performance Analysis.
Our experiments on ImageNet revealed several key findings regarding head architecture (Table 14, Figure 14). Non-linear architectures in the KD head consistently improved performance compared to the linear approach, suggesting that complex architectures better capture teacher predictions. However, increasing complexity in the CE head resulted in performance degradation, most likely due to this head’s focus on limited labeled data, which impaired its capacity to learn generalizable shared feature representations. While dual non-linear heads outperformed their fully linear counterparts, this configuration was less effective than implementing non-linearity solely in the KD head.
When extending our analysis to the other 10 datasets (Table 14, Figure 14), we observed a different pattern. Unlike on ImageNet, the optimal architectural configuration varied considerably across these diverse datasets, with no single configuration emerging as consistently superior. This contrast highlights that the effectiveness of non-linear transformations exhibits strong dataset-specific dependencies and varies according to head functionality. Given these findings, we adopted the linear head architecture for all subsequent experiments, as it achieved comparable performance while providing substantial benefits in computational efficiency and architectural simplicity.
@@
Our analysis under the 16-shot semi-supervised setting revealed intricate relationships between architectural choices and head-wise performance. When employing a non-linear CE branch, we observed a decline in CE head performance from $60.64\%$ to $60.18\%$, despite increased parameters. Conversely, implementing a non-linear architecture for the KD head improved both heads’ performance: CE head accuracy increased from $60.64\%$ to $60.95\%$, and KD head prediction improved from $61.55\%$ to $61.76\%$. However, when applying non-linear architectures to both heads simultaneously, despite enhanced individual performance, the combined performance decreased from $65.97\%$ to $65.59\%$, suggesting that head specialization may compromise joint feature representation quality. These findings underscore the complex interplay between architectural decisions and multi-head learning dynamics in semi-supervised frameworks.
E.3 Additional Dual-Head Analysis

### v2

@@ E.1 Experiments on MobileNet
Figure 10: Results on 10 datasets under few-shot semi-supervision using MobileNetV2 with zero-shot teacher (Radford et al., 2021).
@@
Figure 11: Results on 10 datasets using MobileNetV2 with either zero- or few-shot teacher (Zhang et al., 2021).
@@
To demonstrate the versatility of our DHO approach beyond ResNet (He et al., 2016) and ViT (Dosovitskiy, 2020) architectures, we extended our experiments to the MobileNetV2 (Sandler et al., 2018), which is specifically designed for real-world applications with compact models. We maintained identical experimental settings as described in Section˜C.2, using MobileNetV2 as the student model while distilling from CLIP ResNet50.
@@
As illustrated in Fig.˜11, our DHO consistently outperforms all single-head baseline methods, demonstrating its effectiveness on lightweight model architectures along with ResNet18. Furthermore, Fig.˜11 reveals patterns similar to our ResNet18 experiments regarding few-shot integration. Our method successfully incorporates few-shot teacher knowledge, although we observe that the few-shot teacher does not consistently yield improvements over the zero-shot teacher. Notably, our distilled MobileNetV2 model sometimes achieves superior performance to the zero- and few-shot teachers (ResNet-50) despite having significantly fewer parameters. This pattern of outperforming both zero-shot and few-shot teachers mirrors the observations from our main experiments, further validating the effectiveness of our approach across different architectural families.
E.2 Additional Results with Gradient Surgery Methods
In this section, we present experimental results comparing DHO with PCGrad (Yu et al., 2020b), a well-known gradient surgery method from the multi-task learning literature. PCGrad addresses gradient conflicts by projecting conflicting gradients to resolve conflicts post hoc.
We compare DHO with PCGrad using 8 of the 10 datasets (without Food101, Sun397) with ResNet-18 as the student model. The results are shown in Tab.˜13.
Table 13: Performance comparison of PCGrad with DHO on 8 datasets (average) using ResNet-18.
Method 1-shot 2-shot 4-shot 8-shot 16-shot
CE+KD (logit) 56.2 58.9 61.7 65.8 69.8
PCGrad 56.8 60.3 62.1 67.5 71.6
DHO 60.1 63.7 70.2 75.4 79.1
We observe that PCGrad improves over CE+KD (logit), but still underperforms compared to DHO. Moreover, PCGrad incurs additional memory and computational costs ($\mathcal{O}(|\theta|)$) due to gradient projection and storage, whereas DHO remains lightweight ($\mathcal{O}(d\times C)$) and simple to implement.
The superior performance of DHO can be attributed to its approach of avoiding gradient conflicts at the source by isolating the learning dynamics of each objective via dual heads, rather than resolving conflicts after they arise. Additionally, DHO provides dynamic interpolation capability at inference time between supervised and distillation signals, which standard gradient-based methods do not offer.
E.3 Additional Results with KD Methods
In this section, we present experimental results combining our DHO method with additional distillation approaches: Decoupled Knowledge Distillation (DKD) (Zhao et al., 2022) and Wasserstein Knowledge Distillation (WKD) (Lv et al., 2024). Both DKD and WKD are orthogonal to DHO since their losses can be applied directly to the outputs of $h_{\text{CE}}$. DKD decouples the target class from non-target classes, and WKD computes a kernel matrix within each class. Both approaches require ground-truth labels, restricting their use to the small labeled dataset $\mathcal{D}^{(l)}$.
Despite these limitations, we conducted experiments on ImageNet (Tab.˜14) and on 8 of the 10 datasets (without Food101, Sun397) as shown in Tab.˜15. The results demonstrate that DHO substantially improves the performance of these state-of-the-art KD methods. Notably, WKD+DHO consistently outperforms DHO alone in most settings, demonstrating the extensibility of DHO due to its simplicity.
Table 14: Performance comparison of DKD and WKD with and without DHO on ImageNet. Numbers in parentheses show improvement over the base method.
Method 1-shot 2-shot 4-shot 8-shot 16-shot
DHO 51.8 52.4 52.6 53.3 54.5
DKD 8.9 15.0 20.6 28.2 34.9
DKD+DHO 47.1 (+38.2) 44.3 (+29.3) 40.9 (+20.3) 40.2 (+12.0) 42.4 (+7.5)
WKD 11.0 17.0 17.0 27.7 34.6
WKD+DHO 53.2 (+42.2) 53.3 (+36.3) 53.3 (+36.3) 54.0 (+26.3) 54.8 (+20.2)
Table 15: Performance comparison of DKD and WKD with and without DHO on 8 datasets (average). Numbers in parentheses show improvement over the base method.
Method 1-shot 2-shot 4-shot 8-shot 16-shot
DHO 60.1 63.7 70.2 75.4 79.1
DKD 28.7 41.3 55.4 66.3 73.1
DKD+DHO 46.2 (+17.5) 51.2 (+9.9) 61.8 (+6.4) 69.8 (+3.5) 74.9 (+1.8)
WKD 30.0 38.2 53.3 66.1 73.6
WKD+DHO 59.6 (+29.6) 64.5 (+26.3) 71.2 (+17.9) 75.7 (+9.6) 79.6 (+6.0)
E.4 Additional Results with Adaptive Weighting
To further explore the potential of DHO, we implement an entropy-based adaptive weighting mechanism. Let the entropy of a probability vector $p\in\Delta^{C-1}$ be $H(p)=-\sum_{c=1}^{C}p_{c}\log p_{c}$. We compute the adaptive weight $\alpha$ as:
$\alpha=\frac{\exp(-H(\hat{p}_{\text{CE}}))}{\exp(-H(\hat{p}_{\text{CE}}))+\exp(-H(\hat{p}_{\text{KD}}))}$ (42)
where $\hat{p}_{\text{CE}}$ and $\hat{p}_{\text{KD}}$ are the output probability vectors from $h_{\text{CE}}$ and $h_{\text{KD}}$, respectively. The final prediction is then computed as $\hat{p}=\alpha\cdot\hat{p}_{\text{CE}}+(1-\alpha)\cdot\hat{p}_{\text{KD}}$.
The intuition behind this approach is that lower entropy (higher confidence) predictions should receive higher weights in the final ensemble. When one head produces more confident predictions than the other, the adaptive weighting mechanism automatically emphasizes the more certain prediction. The results show that the entropy-based adaptive weighting method is not proved to be effective in these experiments. While the adaptive weighting does not consistently outperform the fixed interpolation approach, this is likely due to modern neural networks producing overconfident predictions, making entropy an unreliable proxy for uncertainty without proper calibration. However, we believe adaptive weighting could be beneficial with well-calibrated models or alternative uncertainty measures.
Table 16: Performance of ViT-B/16 and ViT-L/14 distilled from ViT-H/14 with entropy adaptive weighting under different percentages of labeled data.
Method ViT-B/16 ViT-L/14
1% 10% 1% 10%
DHO 81.66 82.78 84.59 85.94
DHO+Ent 81.66 82.65 84.60 85.92
E.5 Out-of-Distribution Evaluation upon fully trained model
We provide the evaluation on out-of-distribution datasets with fully trained model in Tab.˜17. DHO significantly outperformed zero-shot baselines on similar-distribution variants (ImageNet-V2, ImageNet-Sketch) across both ViT-B/16 and ViT-L/14 architectures, but showed performance degradation on out-of-distribution datasets (ImageNet-R, ImageNet-A), suggesting increased distribution overfitting from full model training. Interestingly, ViT-B/16 models distilled from ViT-L/14 handled shifted distributions better than those taught by the larger ViT-H/14 DFN (Fang et al., 2023), despite the latter’s superior performance on shifted distributions such as ImageNet-R and ImageNet-A. We attribute this to the shared training background between ViT-B/16 and ViT-L/14 in the CLIP framework (Radford et al., 2021), which appears to better preserve generalization capabilities during the adaptation. This points to an important insight: our method works best on out-of-distributions when the teacher and student models share similar training distributions, suggesting that successful knowledge distillation also depends on the alignment between teacher and student than just the teacher’s raw capabilities.
@@ E.5 Out-of-Distribution Evaluation upon fully trained model
ViT-H (Fang et al., 2023) 632M zero-shot - 83.6 77.2 71.7 92.3 77.4
Appendix F Additional Analysis
F.1 Additional Analysis on Non-Linear Head Design
@@
To further investigate the architectural advantages of dual head optimization, we conducted experiments with non-linear head designs, replacing the linear heads used in our main experiments. We design a non-linear classifier with a sequence of layers: an initial linear projection layer, followed by layer normalization (Ba, 2016), GELU activation (Hendrycks & Gimpel, 2016), dropout (Srivastava et al., 2014), and a final linear classification layer. We compared DHO with three non-linear configurations; DHO+NL-Head-CE: non-linear CE head, DHO+NL-Head-KD: non-linear KD head, and DHO+NL-Head-CE+KD: non-linear both CE and KD heads. All experiments followed the few-shot semi-supervised setting detailed in Section˜C.2.
@@
Table 18: Results of different non-linear head configurations on 11 datasets including ImageNet under few-shot semi-supervision using ResNet-18 with zero-shot teacher (Radford et al., 2021). We report averaged accuracy for 10 visual recognition datasets except the ImageNet.
@@
Figure 12: Results of different non-linear head configurations on 11 datasets including ImageNet under few-shot semi-supervision using ResNet-18 with zero-shot teacher (Radford et al., 2021).
Performance Analysis.
Our experiments revealed key findings regarding head architecture (Tab.˜18, Fig.˜12). On ImageNet, non-linear KD heads consistently outperformed linear ones, suggesting complex architectures better capture teacher predictions. Conversely, non-linear CE heads degraded performance, likely due to overfitting on limited labeled data. While dual non-linear heads outperformed fully linear configurations, they were less effective than non-linearity in the KD head alone.
On the other 10 datasets, optimal configurations varied considerably with no consistently superior approach. This highlights that non-linear transformation effectiveness depends strongly on dataset characteristics and head functionality. Given comparable performance but superior computational efficiency, we adopted linear head architectures for all subsequent experiments.
@@
Analysis under the 16-shot semi-supervised setting revealed complex relationships between architectural choices and head-wise performance as shown in Tab.˜19. Non-linear CE branches decreased CE head performance from $60.64\%$ to $60.18\%$ despite increased parameters. Conversely, non-linear KD heads improved both heads: CE accuracy increased to $60.95\%$ and KD prediction to $61.76\%$. However, dual non-linear heads reduced combined performance from $65.97\%$ to $65.59\%$, suggesting head specialization may compromise joint feature representation. These findings highlight the complex interplay between architectural decisions and multi-head learning dynamics.
F.2 Additional Dual-Head Analysis